OpenStudy (anonymous):

3x - 4y = 7 3x + 2y = -5 When the second equation is subtracted from the first, the resulting equation is _____.

OpenStudy (anonymous):

You subtract each side of the equation. On the left hand side, you get \[ (3x - 4y) - (3x + 2y) = 3x - 4y - 3x - 2y = -6y. \]On the right hand side, you get \[ 7 - (-5) = -12. \]

OpenStudy (anonymous):

12, not -12.

OpenStudy (anonymous):

thank you